At the grocery store, the spring on a scale has a spring constant of 550 N/m. When a bag of apples is hung on the scale it stretches by 4.0 cm.

(a) What is the mass of the apples?  kg
(b) If the scale with the apples in it is pulled down 3.6 cm more and released so that it vibrates up and down

what is the amplitude of vibration? cm
what is the frequency of vibration? Hz
Answer #1

a) at equilibrium

spring force(kx) = mg

where m is mass of apples

=> 550 *0.04=m*9.8

=> m= 2.244 kg

b) time period of oscillation

T = 2*pi*(m/k)^0.5= 2*pi*(2.244/550)^0.5=0.401 sec

SO frequency of vibration f=1/T=2.493 Hz

amplitude of vibration A= 3.6 cm